OpenStudy (anonymous):

A circle has a radius of 1.4 mm and central angle EOG that measures 21 degrees. What is the length of the intercepted arc EG? use 3.14 for pi and round your answer to the nearest tenth. A) 21mm B) 4.4mm C) 1mm D) 0.5mm

zepdrix (zepdrix):

Hey there :)

zepdrix (zepdrix):

If we want to find that piece of arc, what we're really looking for is a `portion of the circumference`. circumference of a circle (distance around the outside) is given by:\[\Large\rm C=2\pi r\]

zepdrix (zepdrix):

But we don't want all of that, we only want 21 out of the full 360 degrees of that arc.

zepdrix (zepdrix):

\[\Large\rm S=\frac{21}{360}\cdot 2\pi r\]

OpenStudy (anonymous):

hiii ok so im doing something wrong because I got a nasty decimal

OpenStudy (anonymous):

im so lost ok do I multiply all of it together or do I need to use parentheses. I don't get it

jimthompson5910 (jim_thompson5910):

as zepdrix shows you would type in (21/360)*2*3.14*1.4 I'm using 3.14 for pi and 1.4 for r

OpenStudy (anonymous):

ok I got 0.5 is that right?

jimthompson5910 (jim_thompson5910):

I'm getting 0.51286666666667 which rounds to 0.5 so it looks good

OpenStudy (anonymous):

yayy great thanks

jimthompson5910 (jim_thompson5910):

yw

zepdrix (zepdrix):

Sorry internet connection went out :(